If five times a number is subtracted from twelve times a number, the result is seven less than twice the number. Find the number
Answer by checkley79(3341)

12X-5X=2X-7
7X-2X=-7
5X=-7
X=-7/5 ANS.
PROOF:
12(-7/5)-5(-7/5)=2(-7/5)-7
-84/5+35/5=-14/5-7
-49/5=-14/5-35/5
-49/5=-49/5